OpenStudy (anonymous):

How do I solve this? Which of the following is a solution of x2 + 5x = -2?

hartnn (hartnn):

solved a quadratic equation before ? know about quadratic formula ?

OpenStudy (anonymous):

Yes and yes :)

hartnn (hartnn):

so, you know you need a,b,c values what did you get them as ?

OpenStudy (anonymous):

a=2 b=5 and c=-2?

hartnn (hartnn):

almost correct! :P but the form is \( ax^2 +bx+c = 0 \) so try again for 'c' value a,b are correct :)

hartnn (hartnn):

actually 'a' is also incorrect

OpenStudy (anonymous):

So would a=1 and c=0?

hartnn (hartnn):

a=1 yes x^2 +5x+2 =0 c= 2 makes sense ?

OpenStudy (anonymous):

Ohh okay I see, yes :)

hartnn (hartnn):

have the formula with you? start plugging in! :)

OpenStudy (anonymous):

Okay so -5 is on the outside and 5^2 -4 (1)(2) over 2 is under the square root sign?

hartnn (hartnn):

go on

hartnn (hartnn):

what is 5^2 - 4(1)(2) =... ?

OpenStudy (anonymous):

Do you multiply the 1 and 2 by -4?

hartnn (hartnn):

yes

OpenStudy (anonymous):

So 25-8?

hartnn (hartnn):

correct

OpenStudy (anonymous):

So now do I find the square root of 17 and divide it by 2 or is that the answer?

hartnn (hartnn):

you're done  ^^ final answer :)